OpenStudy (anonymous):

Multiply and simplify y^2-y-2/y^2-2y+1*y-1/y-2. I am just confused on this one my professor is doing the problem in some way I don't understand, show steps please.

OpenStudy (anonymous):

Is this what you're trying to ask?\[\frac{y^2-y-2}{y^2-2y+1}\times\frac{y-1}{y-2}\]

OpenStudy (anonymous):

\[(y^2-y-2/y^2-2y+1)\times( y-1/y-2)\] factor it. \[(y+1)(y-2)/ (y-1)(y-1)\div (y-1/y-2)\] then cancel. the answer would be. \[y+1/y-1 \]

OpenStudy (anonymous):

yes i dont know how to put the fraction line on the editor

OpenStudy (anonymous):

so basically you have to do group like terms or no.

OpenStudy (anonymous):

you factor the quadratics:\[\frac{(y+1)(y-2)}{(y-1)(y-1)}\times\frac{y-1}{y-2}\]then you cancel the y-1 and the y-2 and get:\[\frac{y+1}{y-1}\]

OpenStudy (anonymous):

to do fractions you use \frac{numerator}{denominator}

OpenStudy (anonymous):

okay I will try that.

OpenStudy (anonymous):

I still can't do the fractions in the equation editor.

OpenStudy (anonymous):

me too. haha how do you that.?

OpenStudy (anonymous):

not for sure

OpenStudy (anonymous):

just type "frac{a}{b}" in the equation editor.